Consider a binomial probability experiment with n=3 and p=0.1. Then, the probability of x=0 is
Answer by stanbon(75887)

Consider a binomial probability experiment with n=3 and p=0.1. Then, the probability of x=0
---------
P(x=0) = 3C0*(0.1)^0(0.9)^3 = 0.9^3 = 0.729
---------------
Cheers,
Stan H.